### v1

@@ 5.3 Confusion at the institutional level
A relevant case law that sanctioned the European Commission [62] demonstrates the importance of taking into account publicly available data to assess the risk of re-identification. The applicant received European funding as a researcher. The funds had been misappropriated, and the costs were ordered to be reimbursed. The Commission published a press release summarizing the decision, without mentioning the applicant’s direct identifiers in order to protect their privacy. However, the researcher brought an action for the annulment of the press release, since it contained identifiable data. The General Court dismissed it, and the applicant appealed its decision to the EUCJ. The Court considered that “information relating to the gender of a person who is the subject of a press release, that person’s nationality, his or her father’s occupation, the amount of the grant for a scientific project and the geographical location of the entity hosting that scientific project, taken together, contain information that may allow the person who is the subject of that press release to be identified, in particular by those working in the same scientific field and familiar with that person’s professional background” and goes on that this circumstance, “ does not allow the risk of identification of the data subject to be regarded as insignificant.”. The judgment avoids discussing anonymization practices in detail but highlights that this issue is subject to misinterpretations, including among public institutions like the Commission. In fact, deleting direct identifiers seems sufficient to achieve anonymization, even by the EU institutions.
5.4 Confusion within firms
@@
Confusion about anonymization practices is also widespread among companies. A good example is the IAB Europe case-law [63]. The company established a set of guidelines aimed at ensuring compliance with the GDPR concerning the collection of browsing data via a TC String (a series of characters coding the user’s preferences). This string could later be used by companies for commercial purposes. The Court ruled that the TC String was a form of personal data, given its capacity to allow individuals to be identified by associating it with additional information (such as an IP address). Despite third parties retaining this additional information, IAB Europe was able to obtain it.
@@
In this regard, the relationship between the EUCJ case-law, employing a risk-based approach to assess the reasonable means likely to be used to re-identify the data, and the WP29 Opinion 5/2014 on anonymization appears complex [90]. On the one hand, discordance persists in the discourse of WP 29 between the zero-risk approach (re-identification must be negligible if not impossible) and the necessity for reasonableness, given that all anonymization techniques are considered imperfect [40]. On the other hand, the need to destroy raw data and to aggregate them in order to achieve anonymization, as required by WP 29, is not met by the case law of the EUCJ, which considers data to be anonymous even if the original data are not deleted, the only relevant aspect being the impossibility (legal rather than technical) to access the additional data that enable re-identification. These contradictions contribute to privacy washing practices by making it difficult to distinguish the company’s bad faith from its lack of knowledge about anonymization methods, especially when data are processed for commercial purposes. In the case of IAB Europe, for example, the company assumed that due to the unavailability of additional information, TC String did not constitute personal data. This mistake is frequently observed among firms. This assertion is supported by numerous decisions adopted by the national DPAs.
@@ 5.5 Confusions arising from the difference between pseudo- and anonymization
One of the most relevant decisions on anonymization dates to 5 September 2024, namely the CEGEDIM SANTE case [23]. CEGEDIM designs and sells secretarial software for the medical sector. The company collected patient health data from doctors who agreed to participate in creating a health data repository. These data were allegedly anonymized with k-anonymity techniques. The decision was based on two key factors: (1) the WP29 test  [^10] Llinkability, re-identification, and inference., which determines whether the data were anonymized, and (2) EUCJ case law. However, the rapporteur isolated a 6-year-old patient with a medical condition, which would suggest pseudonymization, unless re-identification is proved to be “negligible” by “reasonable means”. On this point, the CNIL concluded that the available data could be easily re-identified. However, the company contended that there were few educational materials on anonymization and that the guidelines lacked precision, rendering them unsuitable for legal certainty. Despite CNIL’s rejection of the complaint, the WP 29 Guidelines on anonymization are not updated concerning the current risks associated with k-anonymity. Furthermore, the CNIL has not specified what constitutes a ’negligible risk’. This observation suggests that anonymity-washing may not be a deliberate practice.
Take-away Despite the consistency of the EUCJ case-law relating to personal data, confusion is widespread at the institutional level and among firms as far as the distinction between personal and non-personal data. Such confusion originates from the difficulty in distinguishing anonymization from pseudonymization techniques. This issue could enable unintentional anonymity-washing.
@@ 6 Discussion and overture
Furthermore, in Section 5.1 we show that there is a lack of awareness among practitioners. They often do not recognize that their data can constitute personal data, leading to a failure to implement privacy by design principles [9]. Such misconceptions could be eliminated by adequate training and guidance that would give the right tools to practitioners and engineers to be able to competently asses their datasets and apply anonymization methods.
@@ 6.1 Overture
We believe that one of the most important action to take is to develop a comprehensive anonymization curriculum that could be promoted and distributed by data protection authorities either in the form of training programs offered or thorough and up-to-date educational resources, such as books and hands-on exercises aided with structured guidance on fundamental privacy principles and their real-world applications. Key components should include: (1) A clear explanation of privacy threats and their manifestations in various datasets. (2) An overview of widely accepted privacy definitions (besides k-anonymity and l-diversity, adding differential privacy and cryptographic methods), including their advantages and drawbacks. (3) How these techniques can defend against said privacy threats. (4) Techniques for evaluating privacy technologies and applying them to real-world scenarios. (5) Strategies for auditing privacy risks and implementing mitigation measures in large-scale datasets. (6) Best practices for integrating privacy considerations into broader software engineering projects. (7) Case studies illustrating data breaches and the consequences of inadequate privacy protections. (8) Adding hands-on learning resources, such as Jupyter notebooks and real-world datasets.
Moreover, educational resources should be tailored to diverse audiences. Given the varying levels of expertise among practitioners, privacy education must be designed to accommodate both technical and non-technical professionals. It is also important to emphasize the use of state-of-the-art methods. Practitioners should be trained to critically assess privacy techniques and understand the limitations of traditional methods in modern, large datasets. In order to be sufficiently critical, we should also include adequate privacy risk assessments based on attacker capabilities, data sensitivity, and intended data usage. Finally, we believe that a curriculum of this depth can not be delivered without enhancing the collaboration between academia and industry. We acknowledge that finding a common language between academia and industry is not always straightforward. It takes time and effort: joint initiatives, workshops, and training programs should be encouraged to bridge the gap between theoretical advancements and practical implementation. By realizing these recommendations, organizations and policymakers can move beyond superficial compliance efforts and work toward fostering a robust, meaningful approach to data privacy that will help reduce the prevalence of anonymity-washing and ensure that anonymization practices align with contemporary privacy risks, industrial demands, and capacities.
@@ 6.2 Future work
As a support of the collaboration between academia and industry, we wish to create a repository of guidelines and educational materials using the accumulated knowledge that we have used to write this paper. We envision constructing a proper website [^11] We plan to make a preliminary GitHub repo available upon acceptance of this manuscript. aided with instructions that would help practitioners navigate and find the appropriate guideline, document, or educational resource to their needs.

### v2

@@ 5.2 Case-law on inadequate practices
A relevant case law that sanctioned the European Commission [51] demonstrates the importance of taking into account publicly available data to assess the risk of re-identification. The applicant received European funding as a researcher. The funds had been misappropriated, and the costs were ordered to be reimbursed. The Commission published a press release summarizing the decision, without mentioning the applicant’s direct identifiers to protect their privacy. However, the researcher brought an action for the annulment of the press release, since it contained identifiable data. The General Court dismissed it, and the applicant appealed its decision to the EUCJ. The Court considered that “information relating to the gender of a person who is the subject of a press release, that person’s nationality, his or her father’s occupation, the amount of the grant for a scientific project and the geographical location of the entity hosting that scientific project, taken together, contain information that may allow the person who is the subject of that press release to be identified, in particular by those working in the same scientific field and familiar with that person’s professional background” and goes on that this circumstance, “ does not allow the risk of identification of the data subject to be regarded as insignificant.”. Although the judgment focuses more on privacy protection than anonymization, it highlights the lack of understanding that simply deleting direct identifiers is insufficient to minimize the risk of re-identification. This point is also relevant to anonymization.
@@
Confusion about anonymization practices is also widespread among companies. A good example is the IAB Europe case-law [52]. The company established a set of guidelines aimed at ensuring compliance with the GDPR concerning the collection of browsing data via a TC String (a series of characters coding the user’s preferences). This string, which the company claimed to contain anonymized data, could later be used by companies for commercial purposes. The Court ruled that the TC String was a form of personal data, given its capacity to allow individuals to be identified by associating it with additional information (such as an IP address). Despite third parties retaining this additional information, IAB Europe was able to obtain it.
@@
In this regard, the relationship between the EUCJ case-law, employing a risk-based approach to assess the reasonable means likely to be used to re-identify the data, and the WP 29 Opinion 5/2014 on anonymization appears complex [89]. On the one hand, discordance persists in the discourse of WP 29 between the zero-risk approach (re-identification must be negligible if not impossible) and the necessity for reasonableness, given that all anonymization techniques are considered imperfect [36]. On the other hand, the need to destroy raw data and to aggregate them in order to achieve anonymization, as required by WP 29, is not met by the case law of the EUCJ, which considers data to be anonymous even if the original data are not deleted, the only relevant aspect being the impossibility (legal rather than technical) to access the additional data that enable re-identification. These contradictions contribute to privacy washing practices by making it difficult to distinguish the company’s bad faith from its lack of knowledge about anonymization methods, especially when data are processed for commercial purposes. In the case of IAB Europe, for example, the company assumed that, due to the unavailability of additional information, TC String contained anonymized data. This mistake is frequently observed among firms that find it difficult to distinguish anonymization from pseudonymization and anything in between. This assertion is supported by some decisions adopted by the national DPAs.
@@ 5.3 Confusions arising from the difference between pseudo- and anonymization
One of the most relevant decisions on anonymization dates to 5 September 2024, namely the CEGEDIM SANTE case [26]. It perfectly illustrates the interactions between the WP 29 guidelines on anonymization, the EUCJ’s case-law, and the context-dependent nature of anonymization. Indeed, the nature of the data (health data) and their use (creation of a health data repository) are not neutral. CEGEDIM designs and sells secretarial software for the medical sector. The company collected patient health data from doctors who agreed to participate in creating a health data repository. These data were allegedly anonymized with k-anonymity techniques. The decision was based on two key factors: (1) the WP29 test  [^12] Linkability, re-identification, and inference., which determines whether the data were anonymized, and (2) EUCJ case law. However, the rapporteur isolated a 6-year-old patient with a medical condition, which would suggest pseudonymization, unless re-identification is proved to be “negligible” by “reasonable means”. On this point, the CNIL concluded that the available data could be easily re-identified. However, the company contended that there were few educational materials on anonymization and that the guidelines lacked precision, rendering them unsuitable for legal certainty. Despite CNIL’s rejection of the complaint, the WP 29 Guidelines on anonymization are not updated concerning the current risks associated with k-anonymity. Furthermore, the CNIL has not specified what constitutes a ’negligible risk’. This observation suggests that anonymity-washing may not be a deliberate practice.
Take-away Despite the consistency of the EUCJ case-law relating to personal data, confusion is widespread at the institutional level and among firms as far as the distinction between personal and non-personal data. Such confusion originates from the difficulty in implementing anonymization, which is a probabilistic and context-dependent process, and in distinguishing it from pseudonymization and anything halfway between these two techniques. This issue could enable unintentional anonymity-washing.
@@ 6 Discussion and overture
Furthermore, in Section 5 we show that there is a lack of awareness among practitioners. They often do not recognize that their data can constitute personal data, leading to a failure to implement privacy by design principles [11]. Such misconceptions could be eliminated by adequate training and guidance that would give the right tools to practitioners and engineers to be able to competently assess their datasets and apply anonymization methods.
@@ 6.1 Overture
In order to clarify the complexity of anonymization as a probabilistic and context-dependent process we believe that one of the most important action to take is to develop a comprehensive anonymization curriculum that could be promoted and distributed by data protection authorities either in the form of training programs offered or thorough and up-to-date educational resources, such as books and hands-on exercises aided with structured guidance on fundamental privacy principles and their real-world applications. Key components should include: (1) A clear explanation of privacy threats and their manifestations in various datasets. (2) An overview of widely accepted privacy definitions (besides k-anonymity and l-diversity, adding differential privacy and cryptographic methods), including their advantages and drawbacks. (3) How these techniques can defend against said privacy threats. (4) Techniques for evaluating privacy technologies and applying them to real-world scenarios. (5) Strategies for auditing privacy risks and implementing mitigation measures in large-scale datasets. (6) Best practices for integrating privacy considerations into broader software engineering projects. (7) Case studies illustrating data re-identification and the consequences of inadequate privacy protections. (8) Adding hands-on learning resources, such as Jupyter notebooks and real-world datasets.
Moreover, educational resources should be tailored to diverse audiences. Given the varying levels of expertise among practitioners, privacy education must be designed to accommodate both technical and non-technical professionals. It is also important to emphasize the use of state-of-the-art methods. Practitioners should be trained to critically assess privacy techniques and understand the limitations of traditional methods in modern, large datasets. To be sufficiently critical, we should also include adequate privacy risk assessments based on attacker capabilities, data sensitivity, and intended data usage. Finally, we believe that a curriculum of this depth can not be delivered without enhancing the collaboration between academia and industry. We acknowledge that finding a common language between academia and industry is not always straightforward. It takes time and effort: joint initiatives, workshops, and training programs should be encouraged to bridge the gap between theoretical advancements and practical implementation. By realizing these recommendations, organizations and policymakers can move beyond superficial compliance efforts and work toward fostering a robust, meaningful approach to data privacy that will help reduce the prevalence of anonymity-washing and ensure that anonymization practices align with contemporary privacy risks, industrial demands, and capacities.
We would like to emphasize that we are not suggesting procedural safeguards or checklists, nor do we mean to imply that clear guidance alone can resolve the complexities of anonymisation. Instead, we propose addressing the underlying technical limitations and misunderstandings. Technical guidelines should not be directed at policymakers, but rather at engineers, technical personnel, or even social scientists who need to anonymise data themselves.
@@ 6.2 Future work
As a support of the collaboration between academia and industry, we wish to create a repository of guidelines and educational materials using the accumulated knowledge that we have used to write this paper. We envision constructing a proper website aided with instructions that would help practitioners navigate and find the appropriate guideline, document, or educational resource to their needs.
@@
Acknowledgments.
This work has been partly funded by Project no. 138903, implemented with the support provided by the Ministry of Innovation and Technology from the National Research, Development, and Innovation Fund, financed under the FK_21 funding scheme.
Disclosure of Interests.
The authors declare no competing interests. Project funding is acknowledged in the Acknowledgments section.